OpenStudy (anonymous):

A space shuttle is launched to altitude of 500,000m above the surface of the earth. The shuttle travels at an average rate of 700m/s. How long will it take for the shuttle to reach its orbit?

OpenStudy (anonymous):

Do you have choices? like uh a or b cuz I'm thinkin 500,000 divided by 700 and maybe whatever you get from that is your answer

OpenStudy (mathstudent55):

s = speed; d = distance; t = time s = d/t st = d t = d/s Plug in d and s in the last equation to find the answer.